Given the total cost function; TC = 3Q3 - Q2 + 150Q +50


A.           What is the fixed cost and what does it shows?


B.           What is the variable cost function? What is the variable cost at Q = 2?


C.           What is the maximum output level at the minimum level of average variable cost

D.           At O = 3 what Is the trend of average cost function (is that increasing or decreasing)? Why?


1
Expert's answer
2021-10-01T12:23:39-0400

A.


"FC=50"

B.


"VC=3Q^3-Q^2+150Q"

"VC(Q=2)=320"

C.


"AVC=3Q^2-Q+150"

"\frac{\delta AVC}{\delta Q}=6Q-1"

"Q_{min}=\frac{1}{6}"

D.


"\frac{\delta AVC}{\delta Q}\vert _{Q=3}=17>0"

The trend of average cost function is increasing.